A sphere has an unknown amount of charge distributed uniformly at a distance of 0.56cm from the center of the of the sphere, the electric feld point directly toward centre of the phone which has a magnitude of 1.6x10NC how much charge is on the sphere.
1
Expert's answer
2020-10-15T10:47:32-0400

A uniformly charged sphere creates electric field similar to that of a point charge, therefore, we can use the equation for the field from a point charge:


"E=\frac{1}{4\pi\epsilon_0}\frac{q}{r^2},\\\space\\
q=\frac{r^2}{4\pi\epsilon_0E}=17620\text{ C}."

*In the calculations, we took 0.56 cm = 0.0056 m as distance and 1.6*10=16 N*C as field magnitude.

Since the field lines point toward the sphere, the charge is negative.